# Fisher-Rao Gradient Flow: Geodesic Convexity and Functional Inequalities

### v1

@@
Abstract.
The dynamics of probability density functions has been extensively studied in science and engineering to understand physical phenomena and facilitate algorithmic design. Of particular interest are dynamics that can be formulated as gradient flows of energy functionals under the Wasserstein metric. The development of functional inequalities, such as the log-Sobolev inequality, plays a pivotal role in analyzing the convergence of these dynamics. The goal of this paper is to parallel the success of techniques using functional inequalities, for dynamics that are gradient flows under the Fisher-Rao metric, with various $f$-divergences as energy functionals. Such dynamics take the form of a nonlocal differential equation, for which existing analysis critically relies on using the explicit solution formula in special cases. We provide a comprehensive study on functional inequalities and the relevant geodesic convexity for Fisher-Rao gradient flows under minimal assumptions. A notable feature of the obtained functional inequalities is that they do not depend on the log-concavity or log-Sobolev constants of the target distribution. Consequently, the convergence rate of the dynamics (assuming well-posed) is uniform across general target distributions, making them potentially desirable dynamics for posterior sampling applications in Bayesian inference.
1. Introduction
@@
(1.1) ${\rho^{*}}(\theta)\propto\exp(-V(\theta)),$
in terms of a potential $V(\theta)$. For example, such sampling problems arise in climate science for calibrating climate models [46, 77, 45, 57], engineering for decision making and risk management [87, 26, 36, 14], tissue growth and cell populations models in biological engineering [42, 78, 34], and machine learning for quantifying prediction uncertainties [65, 41, 61, 86, 20]. The demand from these diverse applications has spurred extensive research into efficient and scalable sampling algorithms, a subject explored across applied mathematics, computer science, machine learning, and statistics.
Numerous approaches, based on constructing dynamics of probability density functions, have been proposed in the literature to address the sampling problem. The most widely used examples in Bayesian inference are sequential Monte Carlo (SMC) [29] and Markov chain Monte Carlo (MCMC) [13]. A classical instance in the MCMC category (in continuous time) is the Langevin dynamics [71]
(1.2) ${\rm d}X_{t}=\nabla\log\rho^{*}(X_{t}){\rm d}t+\sqrt{2}{\rm d}B_{t},$
@@
(1.3) $\frac{\partial\rho_{t}}{\partial t}+\nabla\cdot(\rho_{t}\nabla\log\rho^{*})=\Delta\rho_{t},$
which satisfies that when $t\to\infty$, the law of $X_{t}$, namely $\rho_{t}$, converges to $\rho^{*}$, under certain regularity conditions on $\rho^{*}$. The study of the convergence rate has been of significant interest, as it characterizes the mixing time of the dynamics and the efficiency of the sampling algorithm [47, 23].
A remarkable observation is that the Fokker-Planck equation can be formulated as the gradient flow of the Kullback–Leibler (KL) divergence in information theory, known as the Boltzmann relative entropy in kinetic theory [81], with respect to the Wasserstein metric, according to the seminal work of Jordan, Kinderlehrer, and Otto [47]. For this reason, (1.3) is referred to as the Wasserstein gradient flow of the KL divergence. This gradient flow structure has profoundly influenced the analysis of the dynamics [15, 17, 18, 69] and the design of new sampling algorithms [84, 11, 21].
The goal of this paper is to systematically extend the success of the analysis to Fisher-Rao gradient flows which are nonlocal differential equations of great interest in statistics, information geometry and evolutionary game theory. These flows have the potential to achieve fast, universal convergence rates for sampling any target distribution, in contrast to Wasserstein gradient flows. In Section 1.1, we review the results for Wasserstein gradient flows, including functional inequalities (e.g., the log-Sobolev inequality), displacement convexity, and their induced convergence rates of the dynamics. Section 1.2 considers Fisher-Rao gradient flows and explains their significance in sampling applications and the difficulty in analyzing convergence. Section 1.3 presents a summary of our main results for a comprehensive study of geodesic convexity and functional inequalities in the Fisher-Rao setting. We demonstrate the favorable, uniform convergence rates of Fisher-Rao gradient flows across general target distributions. Related works are discussed in Section 1.4, and the organization of this paper is outlined in Section 1.5.
1.1. Wasserstein gradient flows: Functional inequalities and convexity
For the scope of this paper, we adopt the formal approach by Otto [68] to define gradient flows of probability densities. To do so, consider the probability space with smooth positive densities:
(1.4) $\mathcal{P}=\Bigl{\{}\rho\in C^{\infty}(\mathbb{R}^{d}):\int\rho(\theta)\mathrm{d}\theta=1,\,\rho>0\Bigr{\}}\,.$
@@ 1.1. Wasserstein gradient flows: Functional inequalities and convexity
(1.5) $g_{\rho}(\sigma_{1},\sigma_{2})=\langle M(\rho)\sigma_{1},\sigma_{2}\rangle\ \text{for}\ \sigma_{1},\sigma_{2}\in T_{\rho}\mathcal{P},$
[…]
(1.6) $\frac{\partial\rho_{t}}{\partial t}=-{\rm grad}_{M}\mathcal{E}(\rho_{t})=-M(\rho_{t})^{-1}\frac{\delta\mathcal{E}}{\delta\rho}\Bigr{|}_{\rho_{t}},$
@@
(1.7) $\Bigl{\langle}\frac{\delta\mathcal{E}}{\delta\rho},\sigma\Bigr{\rangle}=\lim_{\epsilon\rightarrow 0}\frac{\mathcal{E}(\rho+\epsilon\sigma)-\mathcal{E}(\rho)}{\epsilon},$
@@
(1.8) $\mathcal{E}(\rho)=\text{KL}[\rho\|\rho^{*}]=\int\rho\log\frac{\rho}{\rho^{*}}{\rm d}\theta,$
@@
(1.9) $\frac{\partial\rho_{t}}{\partial t}=-{\rm grad}_{M}\mathcal{E}(\rho)=\nabla\cdot\Bigl{(}\rho_{t}\nabla\log\frac{\rho_{t}}{{\rho^{*}}}\Bigr{)},$
which attains the same form as the Fokker-Planck equation (1.3).
The gradient flow structure of the Fokker-Planck equation naturally leads to Riemannian optimization interpretation for understanding its convergence [69]. In fact, along the gradient flow, it holds that
(1.10) $\frac{\mathrm{d}}{\mathrm{d}t}(\mathcal{E}(\rho_{t})-\mathcal{E}(\rho^{*}))=-g_{\rho_{t}}(\textrm{grad}_{M}\mathcal{E}(\rho_{t}),\textrm{grad}_{M}\mathcal{E}(\rho_{t})).$
If we have the following gradient dominance condition (also known as the Polyak-Łojasiewicz inequality [73]):
(1.11) $\exists\alpha>0,\ g_{\rho}(\textrm{grad}_{M}\mathcal{E}(\rho),\textrm{grad}_{M}\mathcal{E}(\rho))\geq 2\alpha(\mathcal{E}(\rho)-\mathcal{E}({\rho^{*}})),\textrm{ for any }\rho\in\mathcal{P},$
@@
(1.12) $\exists\alpha>0\textrm{ s.t. }\int\rho\bigl{|}\nabla\log\frac{\rho}{{\rho^{*}}}\bigr{|}^{2}\mathrm{d}\theta\geq 2\alpha\int\rho\log\frac{\rho}{{\rho^{*}}}\mathrm{d}\theta,\textrm{ for any }\rho\in\mathcal{P}.$
@@
This is precisely the log-Sobolev inequality [39, 38] for ${\rho^{*}}$ and $f=\sqrt{\frac{\rho}{{\rho^{*}}}}$:
@@
(1.13) $\int\bigl{|}\nabla f(\theta)\bigr{|}^{2}{\rho^{*}}(\theta)\mathrm{d}\theta\geq\frac{\alpha}{2}\int f(\theta)^{2}\log\Bigl{(}\frac{f(\theta)^{2}}{\int f(\theta^{\prime})^{2}{\rho^{*}}(\theta^{\prime})\mathrm{d}\theta^{\prime}}\Bigr{)}{\rho^{*}}(\theta)\mathrm{d}\theta.$
@@
An sufficient condition for the gradient dominance condition is the strong geodesic convexity:
(1.14) $\exists\alpha>0,\ g_{\rho}(\textrm{Hess}_{M}\mathcal{E}(\rho)\sigma,\sigma)\geq\alpha g_{\rho}(\sigma,\sigma),\textrm{ for any }\rho\in\mathcal{P},\sigma\in T_{\rho}\mathcal{P},$
@@
where $\textrm{Hess}_{M}\mathcal{E}(\rho)$ is the Riemannian Hessian of $\mathcal{E}$. For the Wasserstein case $M(\rho)^{-1}\psi=-\nabla\cdot(\rho\nabla\psi)$, the geodesic convexity is also known as the displacement convexity [64]. The Bakry-Emery theorem [8, 9] establishes the connection between the displacement convexity, the log-concavity of $\log\rho^{\star}$, and the Ricci curvature of the sample space, and conditions for general convex Sobolev inequalities involving different relative entropies or divergences were established in [81, 5]. In summary, these functional inequalities and convexity properties play important roles for studying the convergence of dynamics arising from Wasserstein gradient flows [64, 17, 3].
@@
We note that while the above explanations regarding the geometry and gradient flow structures are mostly formal, they can be made rigorous [3, 76]. On the other hand, it is also quite important to use strategically the geometric structure to formally derive the sought inequalities. Then, prove rigorously the resulting inequalities to eventually show the convergence of the dynamics once well-posedness is analysed. In this work, we focus on rigorously proving the functional inequalities.
1.2. Analysis of Fisher-Rao gradient flows
@@
In this paper, our main focus is on the gradient flow dynamics with respect to the Fisher-Rao metric [2, 7], which takes the following form
(1.15) $g_{\rho}^{\mathrm{FR}}(\sigma_{1},\sigma_{2})=\int\frac{\sigma_{1}\sigma_{2}}{\rho}\mathrm{d}\theta,\text{ for }\sigma_{1},\sigma_{2}\in T_{\rho}\mathcal{P}.$
The corresponding metric tensor is $M^{\mathrm{FR}}(\rho)^{-1}\psi=\psi-\mathbb{E}_{\rho}[\psi]$. As an example, the Fisher-Rao gradient flow of the KL-divergence can be expressed as follows:
(1.16) $\frac{\partial\rho_{t}}{\partial t}=-{\rm grad}_{\rm FR}{\rm KL}[\rho_{t}\|{\rho^{*}}]=-\rho_{t}\Bigl{(}\log\rho_{t}-\log{\rho^{*}}-\mathbb{E}_{\rho_{t}}[\log\rho_{t}-\log{\rho^{*}}]\Bigr{)}.$
Recently, convergence of the above Fisher-Rao gradient flow (1.16), namely
(1.17) ${\rm KL}[\rho_{t}\|{\rho^{*}}]=\mathcal{O}(\exp(-t)).$
@@
has been established in [59, 21, 60, 28]. We highlight that the exponential convergence rate here is uniform and independent of $\rho^{*}$. This characteristic stands in sharp contrast to the Wasserstein gradient flow (1.9), where the convergence rate $\mathcal{O}(\exp(-\alpha t))$ depends on the log-Sobolev constant $\alpha$; this constant depends on $\rho^{*}$ and can be very small for anisotropic or multimodal distributions. As a consequence, the uniform convergence rate of the Fisher-Rao gradient flow makes it a potentially desirable dynamics for sampling applications; see [59, 60, 44, 56, 22, 63].
@@
Nevertheless, all the current proofs of the uniform convergence of the Fisher-Rao gradient flow (1.16) rely on the explicit formula of the solution
(1.18) $\rho_{t}(\theta)\propto\rho_{0}(\theta)^{e^{-t}}{\rho^{*}}(\theta)^{1-e^{-t}},$
@@
which is specific to the KL divergence as the energy functional. Moreover, the convergence results require strong assumptions on $\rho_{0}$. This prevents us from extending the convergence framework of Fisher-Rao gradient flows to general energy functionals and initial conditions $\rho_{0}$. On the other hand, the Fisher-Rao gradient flow, although having the form of an ODE at each $\theta$, contains a nonlocal interaction term that conserves the mass: $\mathbb{E}_{\rho_{t}}[\log\rho_{t}-\log\rho^{*}]$. It is non-trivial to analyze such nonlocal terms without using the explicit formula.
@@
In this paper, we address the above challenge by providing a systematic study of functional inequalities and geodesic convexity for the Fisher-Rao gradient flows with general $f$-divergences as energy functionals. Here, the $f$-divergences are defined as
(1.19) $D_{f}[\rho\|{\rho^{*}}]=\int{\rho^{*}}f\Bigl{(}\frac{\rho}{{\rho^{*}}}\Bigr{)}\,\mathrm{d}\theta,$
@@
where $f(1)=0,f\in C^{2}(0,+\infty)$ and $f$ is convex. It is worthy noting that $f$-divergences are known as relative entropies in the context of statistical physics and kinetic theory, see for instance [81, 5, 4].
1.3. Main results

### v2

@@
Abstract
The dynamics of probability density functions have been extensively studied in computational science and engineering to understand physical phenomena and facilitate algorithmic design. Of particular interest are dynamics formulated as gradient flows of energy functionals under the Wasserstein metric. The development of functional inequalities, such as the log-Sobolev inequality, plays a pivotal role in analyzing the convergence of these dynamics. This paper aims to extend the success of functional inequality techniques to dynamics that are gradient flows under the Fisher-Rao metric, with various $f$-divergences serving as energy functionals. Such dynamics take the form of nonlocal differential equations, for which existing analyses critically rely on explicit solution formulas in special cases. We provide a comprehensive study of functional inequalities and the relevant geodesic convexity for Fisher-Rao gradient flows under minimal assumptions. A notable feature of our functional inequalities is their independence from the log-concavity or log-Sobolev constants of the target distribution. Consequently, the convergence rate of the dynamics (assuming well-posedness) remains uniform across general target distributions.
keywords:
Gradient flow, Fisher-Rao metric, Functional inequality, Bayesian inference
{AMS}
62F15, 39B62, 58E10
1 Introduction
@@
(1) ${\rho^{*}}(\theta)\propto\exp(-V(\theta)),$
in terms of a potential $V(\theta)$. For example, such sampling problems arise in climate science for calibrating climate models [isaac2015scalable, schneider2017earth, huang2022iterated, lopez2022training], engineering for decision making and risk management [yuen2010bayesian, cui2016dimension, ghattas2021learning, cao2022bayesian], tissue growth and cell populations models in biological engineering [Hines2014DeterminationOP, simpson2020practical, FCB23], and machine learning for quantifying prediction uncertainties [murphy2012machine, hernandez2015probabilistic, maddox2019simple, yang2021b, chen2021solving]. The demand from these diverse applications has spurred extensive research into efficient sampling algorithms, a subject explored across applied mathematics, computer science, machine learning, and statistics.
Numerous approaches, based on constructing dynamics of probability density functions, have been proposed in the literature to address the sampling problem. The most widely used examples in Bayesian inference are sequential Monte Carlo (SMC) [doucet2009tutorial] and Markov chain Monte Carlo (MCMC) [brooks2011handbook]. A classical instance in the MCMC category (in continuous time) is the Langevin dynamics [pavliotis2014stochastic]
(2) ${\rm d}X_{t}=\nabla\log\rho^{*}(X_{t}){\rm d}t+\sqrt{2}{\rm d}B_{t},$
@@
(3) $\frac{\partial\rho_{t}}{\partial t}+\nabla\cdot(\rho_{t}\nabla\log\rho^{*})=\Delta\rho_{t},$
which satisfies that when $t\to\infty$, the law of $X_{t}$, namely $\rho_{t}$, converges to $\rho^{*}$, under certain regularity conditions on $\rho^{*}$. The study of the convergence rate has been of significant interest, as it characterizes the mixing time of the dynamics and the efficiency of the sampling algorithm [jordan1998variational, chewi2023logconcave].
A remarkable observation is that the Fokker-Planck equation can be formulated as the gradient flow of the Kullback–Leibler (KL) divergence in information theory, known as the Boltzmann relative entropy in kinetic theory [To], with respect to the Wasserstein metric, according to the seminal work of Jordan, Kinderlehrer, and Otto [jordan1998variational]. For this reason, (3) is referred to as the Wasserstein gradient flow of the KL divergence. This gradient flow structure has profoundly influenced the analysis of the dynamics [carrillo2001entropy, CMV03, CMV06, otto2000generalization] and the design of new sampling algorithms [wibisono2018sampling, bernton2018langevin, chen2023sampling].
The goal of this paper is to systematically extend the success of the analysis to Fisher-Rao gradient flows which are nonlocal differential equations of great interest in statistics, information geometry and evolutionary game theory. These flows have the potential to achieve fast, universal convergence rates for any target distribution, in contrast to Wasserstein gradient flows. In Section 1.1, we review the results for Wasserstein gradient flows, including functional inequalities (e.g., the log-Sobolev inequality), displacement convexity, and their induced convergence rates of the dynamics. Section 1.2 introduces Fisher-Rao gradient flows, explaining their significance in sampling applications and the existing methods used to analyze their convergence. Section 1.3 presents a summary of our main results for a comprehensive study of geodesic convexity and functional inequalities in the Fisher-Rao setting. We demonstrate the favorable, uniform convergence rates of Fisher-Rao gradient flows across general target distributions. Related works are discussed in Section 1.4, and the organization of this paper is outlined in Section 1.5.
1.1 Wasserstein gradient flows: Functional inequalities and convexity
For the scope of this paper, we adopt the formal approach by Otto [otto2001geometry] to define gradient flows of probability densities. To do so, consider the probability space with smooth positive densities:
(4) $\mathcal{P}=\Bigl\{\rho\in C^{\infty}(\mathbb{R}^{d}):\int\rho(\theta)\mathrm{d}\theta=1,\,\rho>0\Bigr\}\,.$
@@ 1.1 Wasserstein gradient flows: Functional inequalities and convexity
(5) $g_{\rho}(\sigma_{1},\sigma_{2})=\langle M(\rho)\sigma_{1},\sigma_{2}\rangle\ \text{for}\ \sigma_{1},\sigma_{2}\in T_{\rho}\mathcal{P},$
where $\langle\cdot,\cdot\rangle$ is the duality pairing between the cotangent space $T^{*}_{\rho}\mathcal{P}$ and tangent space $T_{\rho}\mathcal{P}$, and can be identified as the $L^{2}$ inner product given both $\sigma_{1},\sigma_{2}$ are classical functions. Here, $M(\rho):T_{\rho}\mathcal{P}\to T_{\rho}^{*}\mathcal{P}$ is a metric tensor. Given an energy functional $\mathcal{E}:\mathcal{P}\to\mathbb{R}$, the gradient flow under the metric tensor $M(\rho)$ is
(6) $\frac{\partial\rho_{t}}{\partial t}=-{\rm grad}_{M}\mathcal{E}(\rho_{t})=-M(\rho_{t})^{-1}\frac{\delta\mathcal{E}}{\delta\rho}\Bigr|_{\rho_{t}},$
@@
(7) $\Bigl\langle\frac{\delta\mathcal{E}}{\delta\rho},\sigma\Bigr\rangle=\lim_{\epsilon\rightarrow 0}\frac{\mathcal{E}(\rho+\epsilon\sigma)-\mathcal{E}(\rho)}{\epsilon},$
@@
(8) $\mathcal{E}(\rho)=\text{KL}[\rho\|\rho^{*}]=\int\rho\log\frac{\rho}{\rho^{*}}{\rm d}\theta,$
@@
(9) $\frac{\partial\rho_{t}}{\partial t}=-{\rm grad}_{M}\mathcal{E}(\rho)=\nabla\cdot\Bigl(\rho_{t}\nabla\log\frac{\rho_{t}}{{\rho^{*}}}\Bigr),$
which attains the same form as the Fokker-Planck equation (3).
The gradient flow structure of the Fokker-Planck equation naturally leads to Riemannian optimization interpretation for understanding its convergence [otto2000generalization]. In fact, along the gradient flow, it holds that
(10) $\frac{\mathrm{d}}{\mathrm{d}t}(\mathcal{E}(\rho_{t})-\mathcal{E}(\rho^{*}))=-g_{\rho_{t}}(\textrm{grad}_{M}\mathcal{E}(\rho_{t}),\textrm{grad}_{M}\mathcal{E}(\rho_{t})).$
If we have the following gradient dominance condition (also known as the Polyak-Łojasiewicz inequality [polyak1963gradient]):
(11) $\exists\alpha>0,\ g_{\rho}(\textrm{grad}_{M}\mathcal{E}(\rho),\textrm{grad}_{M}\mathcal{E}(\rho))\geq 2\alpha(\mathcal{E}(\rho)-\mathcal{E}({\rho^{*}})),\textrm{ for any }\rho\in\mathcal{P},$
@@
(12) $\exists\alpha>0\textrm{ s.t. }\int\rho\bigl|\nabla\log\frac{\rho}{{\rho^{*}}}\bigr|^{2}\mathrm{d}\theta\geq 2\alpha\int\rho\log\frac{\rho}{{\rho^{*}}}\mathrm{d}\theta,\textrm{ for any }\rho\in\mathcal{P}.$
@@
This is precisely the log-Sobolev inequality [gross1975logarithmic, gross1975hypercontractivity] for $\rho^{*}$} and $f=\sqrt{\frac{\rho}{{\rho^{*}}}}$:
@@
(13) $\int\bigl|\nabla f(\theta)\bigr|^{2}{\rho^{*}}(\theta)\mathrm{d}\theta\geq\frac{\alpha}{2}\int f(\theta)^{2}\log\Bigl(\frac{f(\theta)^{2}}{\int f(\theta^{\prime})^{2}{\rho^{*}}(\theta^{\prime})\mathrm{d}\theta^{\prime}}\Bigr){\rho^{*}}(\theta)\mathrm{d}\theta.$
@@
A sufficient (although not necessary) condition for the gradient dominance condition is the strong geodesic convexity:
(14) $\exists\alpha>0,\ g_{\rho}(\textrm{Hess}_{M}\mathcal{E}(\rho)\sigma,\sigma)\geq\alpha g_{\rho}(\sigma,\sigma),\textrm{ for any }\rho\in\mathcal{P},\sigma\in T_{\rho}\mathcal{P},$
@@
where $\textrm{Hess}_{M}\mathcal{E}(\rho)$ is the Riemannian Hessian of $\mathcal{E}$. For the Wasserstein case where $M(\rho)^{-1}\psi=-\nabla\cdot(\rho\nabla\psi)$, the geodesic convexity is also known as the displacement convexity [mccann1997convexity]. The Bakry-Emery theorem [bakry1985diffusions, bakry2014analysis] establishes the connection between the displacement convexity, the log-concavity of $\rho^{\star}$, and the Ricci curvature of the sample space, and conditions for general convex Sobolev inequalities involving different relative entropies or divergences were established in [To, AMTU]. In summary, these functional inequalities and convexity properties play important roles for studying the convergence of dynamics arising from Wasserstein gradient flows [mccann1997convexity, CMV03, ambrosio2005gradient].
@@
We note that while the above explanations regarding the geometry and gradient flow structures are mostly formal, they can be made rigorous [ambrosio2005gradient, Santambook]. On the other hand, it is also quite important to strategically use the geometric structure to formally derive the desired inequalities and then prove them rigorously. Once well-posedness of the dynamics has been established, these inequalities can be used to establish their convergence. In this work, we focus on rigorously proving the functional inequalities.
1.2 Analysis of Fisher-Rao gradient flows
@@
In this paper, our main focus is on the gradient flow dynamics with respect to the Fisher-Rao metric [rao1945information, amari2016information, ay2017information, laschos2019geometric], which takes the following form
(15) $g_{\rho}^{\mathrm{FR}}(\sigma_{1},\sigma_{2})=\int\frac{\sigma_{1}\sigma_{2}}{\rho}\mathrm{d}\theta,\text{ for }\sigma_{1},\sigma_{2}\in T_{\rho}\mathcal{P}.$
The corresponding metric tensor is $M^{\mathrm{FR}}(\rho)^{-1}\psi=\psi-\mathbb{E}_{\rho}[\psi]$. Historically, this metric was originally referred to as the spherical Hellinger metric, since it arises as the restriction of the Hellinger metric from the space of positive measures to the space of probability measures [hellinger1909neue]. More recently, it has become widely known as the Fisher–Rao metric, which is the terminology we adopt in this work; see LABEL:remark:Fisher-Rao-vs-Hellinger for details.
As an example, the Fisher-Rao gradient flow of the KL-divergence can be expressed as follows:
(16) $\frac{\partial\rho_{t}}{\partial t}=-{\rm grad}_{\rm FR}{\rm KL}[\rho_{t}\|{\rho^{*}}]=-\rho_{t}\Bigl(\log\rho_{t}-\log{\rho^{*}}-\mathbb{E}_{\rho_{t}}[\log\rho_{t}-\log{\rho^{*}}]\Bigr).$
Although this Fisher-Rao gradient flow takes the form of an ODE at each $\theta$, it contains a nonlocal interaction term $\mathbb{E}_{\rho_{t}}[\log\rho_{t}-\log\rho^{*}]$ that conserves the mass. It is nontrivial to analyze such nonlocal terms. Recently, convergence of the above Fisher-Rao gradient flow (16), namely
(17) ${\rm KL}[\rho_{t}\|{\rho^{*}}]=\mathcal{O}(\exp(-t)),$
@@
has been established in [lu2019accelerating, chen2023sampling, lu2022birth, domingo2023explicit]. We highlight that the exponential convergence rate here is uniform and independent of $\rho^{*}$. This characteristic stands in sharp contrast to the Wasserstein gradient flow (9), where the convergence rate $\mathcal{O}(\exp(-\alpha t))$ depends on the log-Sobolev constant $\alpha$; this constant depends on $\rho^{*}$ and can be very small for anisotropic or multimodal distributions. As a consequence, the uniform convergence rate of the Fisher-Rao gradient flow makes it a potentially desirable dynamics for sampling applications; see [lu2019accelerating, lu2022birth, huang2022efficient, lindsey2022ensemble, chen2024efficient, maurais2024sampling].
@@
Nevertheless, most existing proofs of the uniform convergence of the Fisher-Rao gradient flow (16) rely on the explicit formula of the solution
(18) $\rho_{t}(\theta)\propto\rho_{0}(\theta)^{e^{-t}}{\rho^{*}}(\theta)^{1-e^{-t}},$
@@
which is specific to the KL divergence as the energy functional. Moreover, the convergence results require strong assumptions on $\rho_{0}$. This prevents us from extending the convergence framework of Fisher-Rao gradient flows to general energy functionals and initial conditions $\rho_{0}$.
We note that there has also been work using convexity to study the convergence of Fisher-Rao gradient flows. When the energy functional $\mathcal{E}(\rho)$ is strongly relatively convex (in the sense of [lu2018relatively]) with respect to the KL divergence, namely
(19) $\mathcal{E}(\rho^{\prime})\geq\mathcal{E}(\rho)+\int\frac{\delta\mathcal{E}}{\delta\rho}\Bigr|_{\rho}(\rho^{\prime}-\rho)\mathrm{d}\theta+\alpha\rm{KL}[\rho^{\prime}\|\rho],\quad\forall\rho,\,\rho^{\prime},$
for some $\alpha>0$, the Fisher-Rao gradient flow can be shown to converge exponentially fast [yao2024minimizing, Theorem 1] in reverse KL divergence: ${\rm KL}[{\rho^{*}}\|\rho_{t}]\leq e^{-\alpha t}{\rm KL}[{\rho^{*}}\|\rho_{0}]$. In particular, when the energy functional is the KL divergence itself, the strong relative convexity condition (19) holds with $\alpha=1$. Thus, convergence of ${\rm KL}[{\rho^{*}}\|\rho_{t}]$ for (16) is guaranteed; however, convergence in the original KL divergence ${\rm KL}[\rho_{t}\|{\rho^{*}}]$ is not covered in this framework. We note that convergence of the discrete iteration of the Fisher-Rao gradient flow and its connection to mirror descent has also been studied in [chopin2023connection] based on techniques in [aubin2022mirror].
@@
This paper aims to provide a more systematic study of convergence for the Fisher-Rao gradient flows with general $f$-divergences as energy functionals, based on functional inequalities and geodesic convexity that have proven successful for studying Wasserstein gradient flows. Here, the $f$-divergences are defined as
(20) $D_{f}[\rho\|{\rho^{*}}]=\int{\rho^{*}}f\Bigl(\frac{\rho}{{\rho^{*}}}\Bigr)\,\mathrm{d}\theta,$
@@
where $f(1)=0,f\in C^{2}(0,+\infty)$ and $f$ is convex. It is worthy noting that $f$-divergences are known as relative entropies in the context of statistical physics and kinetic theory, see for instance [To, AMTU, HYKE].
1.3 Main results